40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 327 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  504   Thu May 29 16:32:09 2008 steveUpdateSUSetmy oplev is back
I relayed the optics for ETMY-oplev as shown in pictures below.
The reflected beam goes directly to the qpd
Attachment 1: P1020417.png
P1020417.png
Attachment 2: P1020420.png
P1020420.png
  503   Thu May 29 15:58:44 2008 JohnSummaryIOOMC realignment
I repeatedly adjusted the yaw of the upper mirror on the input periscope and re aligned the MC. With the PRM aligned I tried to optimise MC transmission and DC refl simultaneously. I subsequently centred the beams on WFS1/2. Attached is a 30 day trend of MC alignment and transmission.
  502   Wed May 28 14:19:47 2008 steveUpdatePSLkaleidoscope of psl
atm 1: scattering psl table optics from the top of the output periscope f4, 60s @MOPA 3 W
atm 2: scattering psl table optics from the top of the output periscope f4, 20s
atm 3: competing GigE cameras on the north end of psl table
atm 4: yellow "soft" washer to be replaced on psl output periscope
atm 5: ETMY-ISCT in disarray
Attachment 1: pslscat.png
pslscat.png
Attachment 2: pslscat2.png
pslscat2.png
Attachment 3: 2GigEs.png
2GigEs.png
Attachment 4: perwash.png
perwash.png
Attachment 5: etmydisarray.png
etmydisarray.png
  501   Wed May 28 12:51:32 2008 josephbConfigurationComputersTwo more switches mounted
Two more Prosafe 24 port switches have been mounted in the racks, one in 1Y9 and one in 1Y6. (The first one was placed in 1X4).

The one in 1Y9 has been set to an IP address of 131.215.113.251, while the one in 1Y6 is set to 131.215.113.252, and these have been labeled as such.
  500   Tue May 27 16:24:54 2008 tobinConfigurationComputer Scripts / Programsndsproxy
The NDS Proxy is a program that accepts NDS (LIGO Network Data Server) connections from the internet and relays them to
our internal frame-builder, so that you can get DAQ and test-point channel data from off-site.

I stopped the ndsproxy that was running on rana and started it on nodus, its new home. This will be
documented in the wiki.

So far I haven't found a mechanism by which the ndsproxy was restarted automatically on rana. Has it just been
restarted by hand?

The ndsproxy stuff lives in target/ndsproxy. Restarting it seems to be just a matter of running "start_ndsproxy" in
that directory.
  499   Sun May 25 22:33:00 2008 ranaUpdateSUSETMY Oplev Work
I found the ETMY table in disarray and put the panels back on and put the ETMY OL laser back on the quad.

The next thing to do is clean up the table (there's a lot of junk on it) and then put in a lens within
6" of the laser to focus the beam on the quad (no metal diving boards and the lens should be either
uncoated (from our Edmunds collection) or a red lens, not 1064).

Then we have to put the beam cover back on between the viewport and the table.
  498   Sun May 25 21:14:14 2008 tobinConfigurationComputersEPICS proxy server
I set up an EPICS gateway server on Nodus so that we can look at 40m MEDM screens from off-site.
The gateway is set up to allow read access to all channels and write access to none of them.

The executable is /cvs/cds/epics/extensions/gateway; it was already installed. A script to start
up the gateway is in target/epics-gateway. For the time being, I haven't set it up to start itself
on boot or anything like that.

To make it work, you have to set the environment variable EPICS_CA_ADDR_LIST to the IP address of
Nodus. For instance, something like this should work:

setenv EPICS_CA_ADDR_LIST 131.215.115.52

On Windows you can set up environment variables in the "System" Control Panel. On one of the tabs
there's a button that lets you set up environment variables that will be visible to all programs.

On Andrey's machine I installed the Windows EPICS extensions, i.e. MEDM and its friends. I also
installed the cool Tortoise SVN client which lets you interact with SVN repositories through
the windows explorer shell. (The right-click menu now contains SVN options.) I checked out
the MEDM directory from the 40m SVN onto the desktop. You should be able to just right-click in
that window and choose "SVN Update" to get all the newest screens that have been contributed to
SVN; however, there are currently some problems with the 40m SVN that make that not go smoothly.

At the moment on Andrey's (Windows) machine you can go into the MEDM folder and double-click on
any screen and it will just work, with the exception that not all the screens are installed
due to SVN difficulties.
  497   Sun May 25 20:30:25 2008 ranaSummaryPSLPMC Mode Matching
I checked the PMC mode matching by ramping the gain down to -10 dB (from +20 dB) and
moving the DC offset around until it caught lock on the different HOMs. Then I recorded
the output power (PMCTRANSPD). The DC offset on this EPICS channel was -0.013 V, so I
used its AOFF field to zero this out. Here is a list of the power in the largest modes:
Mode    Power (V)
----    ---------
00        2.7
10        0.2
04        0.04
02        0.02
BE        0.36      **Bull's Eye mode is TEM02 + TEM20. This can be fixed by lens adjustment.


N.B. To make a PNG file with DTT, just make an EPS file -- then use the eps2png perl script.
Attachment 1: pmc.png
pmc.png
  496   Sun May 25 19:33:16 2008 ranaUpdateIOOMC Bad after Re-alignment
The MC pointing was off and the transmitted power was down after John and Andrey brought it back after the bootfest.

I tried getting it back on Friday but was unsuccessful. Today, after the Phoenix Landing, I got it back to someplace
reasonable, but it still seems to be far off. I will check with Rob before we recenter and of the QPDs.

I had to move all of the MC SUS and also align the beam using the IOO periscope. The attached PDF shows some trends
over the last 80 days. You can see that the drop in MC TRANS is about the same as the drop in PMC TRANS.
Attachment 1: e.pdf
e.pdf
Attachment 2: pmc.png
pmc.png
  495   Sun May 25 16:20:27 2008 ranaConfigurationComputersjoinPDF
I have installed joinPDF 2.1 on rosalba. Since its written in Java, I didn't have to tinker with it at all to work on a 64-bit machine. Now Caryn can put all of her plots into 1 file.
  494   Fri May 23 21:21:52 2008 CarynSummaryGeneralfiltering mode cleaner with wiener filter
I tried filtering some saved MC_L data (from Mon May19 4:30pm) with multiple MISO filters of different orders, with various sampling rates, at different times. Plotted the max rms error (where error is signal minus signal-estimate). 2min of data (around Mon May19 4:30pm) were used to calculate each filter. And each filter was applied to data at later times to see how well it performed as time progressed. Plots are attached. There appears to have been a disturbance during the 3rd hour. Rana pointed out perhaps it would be better to use data from the evening rather than during the day.
Attachment 1: error_vs_N_for_different_times_64Hz.pdf
error_vs_N_for_different_times_64Hz.pdf
Attachment 2: error_vs_N_for_different_times_128Hz.pdf
error_vs_N_for_different_times_128Hz.pdf
Attachment 3: error_vs_N_for_different_times_256Hz.pdf
error_vs_N_for_different_times_256Hz.pdf
Attachment 4: error_vs_N_for_different_times_512Hz.pdf
error_vs_N_for_different_times_512Hz.pdf
Attachment 5: error_vs_srate_for_different_times_256.pdf
error_vs_srate_for_different_times_256.pdf
Attachment 6: error_vs_srate_for_different_times_512.pdf
error_vs_srate_for_different_times_512.pdf
Attachment 7: error_vs_srate_for_different_times_1024.pdf
error_vs_srate_for_different_times_1024.pdf
Attachment 8: error_vs_time_for_different_srates_256.pdf
error_vs_time_for_different_srates_256.pdf
Attachment 9: error_vs_time_for_different_srates_512.pdf
error_vs_time_for_different_srates_512.pdf
Attachment 10: error_vs_time_for_different_srates_1024.pdf
error_vs_time_for_different_srates_1024.pdf
  493   Fri May 23 08:24:24 2008 ranaAoGTreasureYoichi Aso has arrived !
Attachment 1: Colossus_back_(800_x_600)-thumb.jpg
Colossus_back_(800_x_600)-thumb.jpg
  492   Thu May 22 11:25:19 2008 josephbConfigurationComputers 
One of the new Netgear Prosafe 24 port switches was mounted in the 1X4 rack,, roughly in the middle, away from the top and bottom rack mounted electronics. At the moment, its IP has been set to 131.215.113.250, gateway 131.215.113.2 (which is what I saw as the only listed gateway on linux1 using route -n) and mask 255.255.255.0.

I'm planning to set the next three IP address for the switches as *.251, *.252 and *.253, which don't look to have been used yet.
  491   Thu May 22 11:21:45 2008 steveBureaucracySAFETYearly surf sudent
Caltech under grad Eric Mintun received the 40m safety training.
Now he has to read and sign SOP of laser and ifo.
He'll be working with GigE cameras with Joe
  490   Wed May 21 15:21:33 2008 robUpdateComputer Scripts / Programsautolockers and cron

I added hourly cron jobs to op340m to ensure that

MC autolocker
FSS Slow Servo
PSL watch

are running. I've also edited the wiki procedure to reflect the fact that these no longer need to be restarted by hand.
  489   Tue May 20 18:33:01 2008 Andrey, JohnConfigurationIOOMode Cleaner is locked again

It was noticed by Mr.Adhikari earlier today that the MC became unlocked at about 11AM.

There is no clear understanding what caused the problem.

Trying to restore the modecleaner locking, we noticed with John that the beam was not centered on the wavesensors (WFS1. WFS2 on the screen "C1IOO_LockMC.adl"). We decided to adjust the beam position moving slightly the bias sliders for pitch and yaw degrees of freedom for MC1.
This allowed to make the MC locked.

Old positions for the MC1 sliders: Pitch = 2.9934, Yaw = -0.6168;
New positions --------//---------: Pitch = 3.0604, Yaw = -0.7258.

At the same time, FSS for PSL is still showing the values in the range 0.720 - 0.750 which is lower than the usual values. The indicator for FSS value is yellow when it is below 0.750.
  488   Tue May 20 09:28:42 2008 steveBureaucracySAFETYsafety traning for 40m
Tara Chalernsongsak, new gradstudent for K. Libbrecht was introduced to the basics of the 40m operations.
  487   Mon May 19 11:49:57 2008 steveUpdateSUSetmy oplev laser replaced
The 9 years old Uniphase HeNe was replaced by a new JDSU 1103P
Power output ~3.5 mW, reflected light back on qpd ~140 microW, 12,800 counts
Andrey will get the ~2.5 mm od spot on the qpd smaller by replacing
the f1000 lens
  486   Sun May 18 18:59:15 2008 ranaConfigurationComputerscron and hosts
I added rosalba to the hosts file for the control room machines (131.215.113.103).

I also removed the updateddb cron from our op440m crontab because it was running at 5 PM
even though I had set it to run at 5:57 AM. If it still runs then, it must be because of
another crontab.
  485   Sun May 18 18:44:48 2008 ranaSummarySUSOptical Lever SUM Trend - 80 days
I used the OL-Trend.xml dataviewer template to make this plot. Looks like the ETMY optical lever
slowly degraded over the last few months and then finally died 10 days ago. Would someone please
replace this laser and tune the lens position to minimize the spot size on the quad?
Attachment 1: e.pdf
e.pdf
  484   Sun May 18 18:14:30 2008 ranaSummaryComputer Scripts / Programsautlockers and cron
Today I noticed that the MC was unlocked, the MC autolocker was off, the PSL autolocker was off,
and the PSL Slow loop was off.

Reading through .log files and our elog it seems like things have been this way since Tuesday
when Andrey went around rebooting computers to bring things back.

And it looks like the ETMY optical lever is dead.

I restarted the PSL & MC scripts on op340m. I also locked and aligned the X arm to verify that
not everything is broken. The Y Arm is unalignable until we replace the HeNe.
  483   Fri May 16 17:27:55 2008 AndreyOmnistructureGeneralToilets are broken, do not use them !!!

Both toilets in 40-meter were constantly flushing, the leaking water was on the floor inside of the restrooms, so

BOTH RESTROOMS ARE CLOSED TILL MONDAY


I have heard the constant loud sound of flushing water, opened the door, and was unpleasantly surprised because all the floor was under the layer of water and the toilets were constantly flushing. I called security at X5000, a plumber came in and told that a team of plumbers needs to repair the flushing system after the weekend. The plumber today just shut off the flushing water, wiped off the floor and told not to use the restrooms in the weekend. We should expect a team of plumbers on Monday.

Sinks are working, so you can wash your hands.
  482   Fri May 16 14:38:50 2008 josephbSummaryCamerasTwo cameras setup
I've changed the pickoff setup from yesterday for the GigE cameras to include a 33% beam splitter (first one I could find). The reflection is going to the GC650 (CCD camera) while the transimission is going to the GC750 CMOS camera. This means the CMOS camera has roughly twice the light incident as the GC650 and should be kept in mind in all comparisons. The distances from the beam splitter are approximately the same both cameras, but some more accurate positioning might be useful.

Its very easy to get the GC650 camera into a bad state where you need to go out and cycle the power (simply unplug and re-plug in the power supply either at the camera or outlet). If the ListCamera program doesn't see it, this is probably necessary.

Andrey added at 6.30PM: Actually the 650 camera keeps crashing constantly. Every time I attempt to capture an image, the camera fails.
  481   Thu May 15 16:24:18 2008 josephbSummaryCameras 
The GC750 camera is currently looking at a very small pickoff of the PSL output (transmission of a Y1-1037-45-S mirror). The plan is to take images tomorrow with it and the GC650 from the same spot and do comparisons.

For those interested, the camera can be run with two codes, from mafalda. Use ssh -X mafalda to login, to allow the live stream to work with the SampleViewer code. The codes can be found in:

/cvs/cds/caltech/target/Prosilica/40mCode/Snap

and

/cvs/cds/caltech/target/Prosilica/bin-pc/x86/SampleViewer

Type Snap --help for a list of options for that program. Click the circle looking thing in SampleViewer to start the live stream. Note only 1 of the two programs can be running at a time, and the only way to change settings (such as exposure length) is with Snap at the moment.
  480   Thu May 15 14:39:33 2008 CarynSummaryPEMfiltering mode cleaner with mic
Tried filtering for mode cleaner data(C1:IOO-MC_L) using a siso-firwiener filter and microphone data(C1:PEM-AS_MIC) for noise input. The noise reduction in mode cleaner data using the microphone-filter is comparable to the noise reduction when an accelerometer(C1:PEM-ACC_MC1_X) filter is used. See attached graphs.
Attachment 1: MC_L_with_PEM-AS_MIC_filter.pdf
MC_L_with_PEM-AS_MIC_filter.pdf
Attachment 2: MC_L_with_PEM-ACC_MC1_X_filter.pdf
MC_L_with_PEM-ACC_MC1_X_filter.pdf
  479   Thu May 15 12:05:49 2008 josephbConfigurationPSLPath to PSL Position QPD
The 50/50 beamsplitter that was being used as the last turning mirror to the PSL Position QPD has been replaced with a Y1-1037-45-S plate. This turning mirror was also moved 4" farther along the beam path, so as to produce as small (few microwatts) transmission through the plate. The lensing optics were also shifted so as to maintain a focused beam on the photodiode. Lastly, the rotating ND filter was increased from 1.5 to 2.0 to reduce the incident power on the photodiode, since twice the power is now reaching it.

The small beam on transmission will be used by the digital cameras as a test beam.
  478   Thu May 15 10:40:21 2008 steveHowToGeneralLisa Goggin, PhD
Lisa Goggin successfully defended her thesis on May, 13 2008

"A Search For Gravitational Waves from Perturbed Black Hole Ringdowns in Ligo Data"

She started out as a surf student in the 40m.

Congratulation!
Attachment 1: lisa.JPG
lisa.JPG
  477   Wed May 14 14:05:40 2008 AndreyUpdateComputersComputer Linux-2, MEDM screen "Watchdogs"

Computer "Linux-2", MEDM screen "C1SUS_Watchdogs.adl": there is no indication for ETMY watchdogs, everything is white. There is information on that screen "C1SUS_Watchdogs.adl" about all other systems (MC, ETMX,...), but something is wrong with indicators for ETMY on that particular control computer.
  476   Wed May 14 13:14:19 2008 AndreySummaryComputersReflective Memory Network is restored

Reflective Memory Network is restored, all watchdogs and oplevs are returned to the "enabled" state.

In order to revive the computers, several things were done.

1) Following Mr. Adhikari's elog entry #353, I walked around the interferometer room, and switched off the power keys in all crates with computers whose names are contained in the MEDM Reflective Memory screen, including the rack with the framebuilder. By the way, it was nontrivial to find the switch in the 1Y4 crate that would shut off/on processors "c1susvme1" and "c1susvme2": the switch turned out to be located at the rear side of the crate, and it is not a key but it is a button.

2) I was trying to follow wiki-40 computer restart procedures, but every time that I was trying to run "startup.cmd" screen from the corresponding target subdirectory, I got the error message "Device or resource busy".
By the way, one more thing was learned: if you firstly open in terminal burtgooey, select the snap file, then reboot the processor, and then will try to burt-restore it, you will get the message "Status Not OK". In order to really burt-restore the processor which was recently rebooted, you need to close the terminal with burtgooey and open burtgooey in a new terminal window which should be opened after rebooting the processor.

Feeling that my activities according to wiki-40 procedures do not revive computers, I invited Alex Ivanov.

3) Alex tried to touch the memory card in "c1iovme" in rack 1Y2, because once before this card failed causing network problems, but this did not help.

4) We shutted off and restarted again (pressing the power-switching button) the black Linux machine "c1dcuepics" (located in the very bottom below the framebuilder). Alex says that this machine is responsible for all EPICS. It was not restarted for 182 days, and probably some process there went wrong.

After restarting this machine "c1dcuepics" we were able to follow wiki-40 procedures for restarting all other computers (whose names are on the MEDM RFM network). We ran correcponding "startup.cmd" files and burt-restored them without error messages.

Now all the computers work and communicate in a proper way.

Mr. Joseph Betzwiezer was helping me with all these activities (we decided that it is more important that cameras for now), thanks to him. But our joint skills turned out to be insufficient, so Alex Ivanov's contribution was the most important.
  475   Tue May 13 10:38:28 2008 steveUpdateComputersrfm network is down
The RFM network went down yesterday around 5pm
Only c1susvme1 is alive but it's timing is off.
Andrey is bringing the network up.

Andrey wants to make an addition first that our situation is very much similar to that described by Rana in his elog entry # 353 (March 03). All of the rectangular boxes are red, except for SUS1-c1susvme1 and AWG (only these two rectangles are green).
  474   Tue May 13 10:15:52 2008 steveUpdateSUSrestored damping of BS & PRM
I think our janitor was cleaning too heavy handedly.
The BS and PRM sus damping were lost.
They were restored.
  473   Fri May 9 10:15:36 2008 josephbUpdateComputersNodus has moved
Steve and myself moved Nodus from under the table in the control room, to just above the Rana computer in the control room rack.
  472   Fri May 9 08:40:24 2008 steveUpdateSUSETMY sus damping restored
ETMY lost damping at 19:10 last night.
There was no seismic event than.
Sus damping was restored this morning.
  471   Thu May 8 16:40:36 2008 josephbConfigurationCamerasGige Camera currently on PSL table
Andrey and myself were working on the PSL table today, using a pickoff of a pickoff of the main beam (adding a microscope slide to pickoff ~4% of the original pickoff) to the GC750 GigeCam.

At the time we left, we scanned the area with a beam scan and didn't see any new stray beams, and nothing in any useful beam paths should have changed. We also strung a Cat 6 cable from the control room switch out to the PSL table in the cable trays, and then above the PSL table.

Currently, its not as well aligned as it could be, and also requires a very low exposure setting, of -E 50 or so to avoid saturation.
  470   Thu May 8 02:06:13 2008 ranaSummaryCOCThermal Lensing in the ITMs and BS may be a problem
The iLIGO interferometers start to see thermal lensing effects with ~2W into the MC, a recycling
gain of ~50, and a beam waist on the ITMs of ~3.5 cm.

At the 40m, the laser power into the MC is 1/2 as much, the recycling gain is 4-5x less, but the
beam on the ITM has a 3 mm waist. So the power in the ITM bulk is 10x less but the power density
is 100x more
. Seems like the induced lens in the ITM bulk might be larger and that if there's
significant absorption on the ITM face (remember our Finesse is 4-5x higher) the beam size in the
arm cavity may also change enough to measure.

Someone (like Andrey) should calculate how much the beam sizes change with absorbed power.
  469   Thu May 8 01:50:25 2008 ranaSummaryASCArm Cavity HOM Resonances
Nothing new, but I calculated the frequencies of the first 22 higher order transverse modes and thought I might as well list them here.

To do this I took formula (23) from page 762 of Siegmans book and put it into this form:
         f_fsr
dfmn =   ----- * (m+n) * acos(sqrt(g1*g2))
           pi

and then calculated them from m+n = 1..22 (22 is not a magic number).

I also used the 'mod' function of matlab to calculate the frequency mod FSR so that we would know how far away
from a cavity resonance it is. I took as parameters: Larm = 38.55 m, Ritm = 1e6 m, Retm = 57.1 m. Kirk measured
the arm length some time ago; we need to measure the arm g-factor...maybe we'll put Tobin on this when he comes
by for a visit.

1.1936 (TEM01, TEM10)
2.3871
3.5807
0.8859 (TEM22, TEM13, TEM31)
2.0795
3.2730
0.5782
1.7718
2.9654
0.2706 (TEM55, ...)
1.4641
2.6577
3.8512
1.1564
2.3500
3.5436
0.8488
2.0423
3.2359
0.5411
1.7347
2.9282
  468   Thu May 8 01:07:24 2008 ranaSummaryLSCFrequency Noise test: MC Trans Backscatter
There is a wandering hump in the MC_F spectrum. It can move around on the time
scale of seconds between 40 and 200 Hz. It has an amplitude ~5-50x above the background spectrum. This seems new; I don't remember it
from a year ago. It is there in the IFO unlocked as well as the IFO locked as well as the locked + CM mode.

Tapping the AS table and/or the PSL table enclosures produces a broadband increase in the MC_F spectrum but doesn't
selectively effect the hump.

We thought it might be backscatter from the MC TRANS path and so we stuck in one of Steve's cool black glass V's into
this space. No effect. We should design a black glass V dump which we can replicate in large quantities for us and for
the sites. Something like the one on the LSC PDs, but with a 1 sq. inch opening area and a 2 inch depth.


We have also done this on the MC2 - TRANS beam before. No noise reduced there either.

The noise hump is appearing in MC_F but not in CARM_IN1 (after the CM handoff) so it seems like the MC has enough gain
to squash it. This also exonerates the MC REFL path since anything there would not be effected by the MC servo gain and
so would be visible in CARM.

My best guess is that there is something really, really scattery on the PSL table. But for now it looks like this is not a
big factor in the locking
issues.
  467   Wed May 7 15:25:41 2008 robConfigurationLSCAP33 -> POX33

Quote:

I am in the process of switching the POX166 and AP33 photodetectors, so that they become POX33 and AP166. The IFO_CONFIGURE buttons won't work until I finish.


Done. We're now in the 40m CDD configuration.
  466   Tue May 6 17:28:39 2008 robConfigurationLSCAP33 -> POX33

I am in the process of switching the POX166 and AP33 photodetectors, so that they become POX33 and AP166. The IFO_CONFIGURE buttons won't work until I finish.
  465   Mon May 5 15:53:14 2008 steveUpdateVACdrypump replaced & readbacks are out
The small turbo pump of the annulus system is tp3
It's drypump was replaced at 1.4 Torr after 8 months of operation.
The rebuilding cost of this SH100 drypump is getting ridiculously high $1349.
I'll look for a replacement.

Joseph rebooted c1vac2 card on Apr 24 entry#441
This restarted the readbacks of the turbos and ion pumps for a while.
That was nice. The readbacks are not working again

  464   Mon May 5 11:04:30 2008 robOmnistructureComputersNetwork setup

Mafalda was not connected to the network, and so our DMF-based seisBLRMS has not been running for ~1 week. I traced this to a broken ethernet cable connecting mafalda to the network switch in the rack next to the B&W printer. This cable has a broken connector at the switch side, which means it can't stay connected if there's any tension. It needs to be replaced.
  463   Thu May 1 12:46:02 2008 josephbConfigurationComputersNodus gateway is up
The computer Nodus is now acting as a gateway machine between the GC network and the martian network in the 40m. It has the same passwords as the rana gateway machine.

Its name on the GC side is nodus (ip: 131.215.115.52) and on the martian side is nodus113 (ip: 131.215.113.200). Will need to update the hosts file on the control room machines so you can just use the name nodus113 rather than the full ip.

Software is still being added to the computer, and it will remain in parallel with the rana gateway machine until everything has been working properly for a week or so.
  462   Thu May 1 08:31:51 2008 steveUpdateSUSearthquake trips etmy & mc1
Earthquake at Lake Isabel magnitude 4.4 at 1am today shakes the 40m
ETMY and MC1 watchdogs tripped.
Sus damping restored.
  461   Wed Apr 30 20:48:58 2008 AndreySummaryPEMNew Weather Channels

I created the new channels for the weather station, all letters are capital ones. They are of the form "C1 : PEM-WS_PARAMETER" where "PARAMETER" is temperature, pressure, wind,... characteristics (names are self-obvious).

These new weather channels are indicated on the "Weather Checklist" MEDM screen. Also, units of pressure were changed from Pascal to torr and mbars.

The new weather channels are also visible in Dataviewer. I updated the template, and as an example of Dataviewer data I attach the following 5-hour trends of weather parameters from 3.30PM to 8.30PM on April 30th.
Attachment 1: April30-5hours.png
April30-5hours.png
  460   Tue Apr 29 21:30:49 2008 AndreyUpdatePEMIn the process of renaming channels for Weather Station

I startted renaming channels for the weather station, and I will continue this tomorrow, on Wednesday.

I have restarted 'c1pem1' several times and reconfigured "C0DCU1" on the framebuilder MEDM screen.

Framebuilder now does not work.
  459   Tue Apr 29 21:09:12 2008 ranaDAQCDSFE Filters
These are new FE filters for downsampling and upsampling. We will be going from native hardware sampling rates of 64k down to 32k, 16k, and 2k.

The attached plot shows these filters. They are 3dB ripple, 40 dB stopband, 4th order elliptic filters in which I have moved the zeros around
into good places (e.g. to the Nyquist frequency).

I'm also attaching the .txt file containg the filter coefficients and the design strings. The filters are called x2, x4, and x32, for the
D2, D4, and D32 downsampling, respectively.
Attachment 1: fefilters.jpg
fefilters.jpg
Attachment 2: fefilters.txt
# FILTERS FOR ONLINE SYSTEM
#
# Computer generated file: DO NOT EDIT
#
# MODULES ULYAW
#
################################################################################
### ULYAW                                                                    ###
################################################################################
# SAMPLING ULYAW 65536
... 28 more lines ...
  458   Mon Apr 28 23:44:33 2008 AndreyUpdateComputer Scripts / ProgramsWeather.db

I was trying to figure out how to modify the file "Weather.db" so that the atm.pressure would be recalculated from Pa to bar before appearing in the EPICS screen, but so far I did not succeed. I restarted processor "c1pem1" several times. I will continue this tomorrow, and also I will modify the nmaes of the weather channels.
  457   Sun Apr 27 22:57:15 2008 ajwDAQComputersbr40m?

Quote:

The testpoint manager (which runs on fb40m) crashed this afternoon. Upon re-starting it, I found there was a rogue dtt process on op440m and also a daqd daemon running on br40m. One or both of these caused the tpman to crash. br40m is the frame broadcaster, which is never used here as we don't run DMT. I killed the daqd process there.

The way to find if there is a rogue process is to watch the output to the console from the tpman when you start it:

Allocate new TP handle 56 by 131.215.113.203
Allocate new TP handle 57 by 131.215.113.203
Allocate new TP handle 58 by 131.215.113.203
Allocate new TP handle 59 by 131.215.113.203
Allocate new TP handle 60 by 131.215.113.203
Allocate new TP handle 61 by 131.215.113.203
Allocate new TP handle 62 by 131.215.113.203
Allocate new TP handle 63 by 131.215.113.203
Allocate new TP handle 64 by 131.215.113.203
Allocate new TP handle 65 by 131.215.113.203
Allocate new TP handle 66 by 131.215.113.203
Allocate new TP handle 67 by 131.215.113.203
Allocate new TP handle 68 by 131.215.113.203


If you see something like this, with a new TP handle being allocated every few seconds, you need to log in to the corresponding host and kill whatever process has run away.


I *think* Alex is responsible for the daqd daemon running on br40m (he set up some new stuff recently, a data concentrator and broadcaster); I'll make sure he sees this post.
  456   Sun Apr 27 18:11:58 2008 robDAQComputersbr40m?

The testpoint manager (which runs on fb40m) crashed this afternoon. Upon re-starting it, I found there was a rogue dtt process on op440m and also a daqd daemon running on br40m. One or both of these caused the tpman to crash. br40m is the frame broadcaster, which is never used here as we don't run DMT. I killed the daqd process there.

The way to find if there is a rogue process is to watch the output to the console from the tpman when you start it:

Allocate new TP handle 56 by 131.215.113.203
Allocate new TP handle 57 by 131.215.113.203
Allocate new TP handle 58 by 131.215.113.203
Allocate new TP handle 59 by 131.215.113.203
Allocate new TP handle 60 by 131.215.113.203
Allocate new TP handle 61 by 131.215.113.203
Allocate new TP handle 62 by 131.215.113.203
Allocate new TP handle 63 by 131.215.113.203
Allocate new TP handle 64 by 131.215.113.203
Allocate new TP handle 65 by 131.215.113.203
Allocate new TP handle 66 by 131.215.113.203
Allocate new TP handle 67 by 131.215.113.203
Allocate new TP handle 68 by 131.215.113.203


If you see something like this, with a new TP handle being allocated every few seconds, you need to log in to the corresponding host and kill whatever process has run away.
  455   Sun Apr 27 05:09:30 2008 ranaConfigurationIOOMC WFS Whitening turned on
I hardwired on the MC WFS whitening filters.

The MAX333A switches which choose between whitening and bypass on that board were in the bypass position
because the Xycom220 connections are not there. So the control switch gets +15V but there is no pull
down to set it to the whitened mode.

The least invasive (easiest) change I could do was to tie all of those inputs to ground. This pulls a few mA
through the pull-down resistors but is otherwise innocuous. All of these control lines come in on the A-row
of the P1 connector, so I was able to solder a single wire across all of them to ground them all.

The WFS2 board had a blown electrolytic capacitor on the -15 V line and so there was probably some extra noise
getting in that way. I couldn't find any extra SMD to replace it so I cut the legs off of a 22 uF polarized
tantalum and stuck it in there. Its even close to being the same color. I checked out the other caps, and they were all
close to 68 uF as spec'd. This one had luckily blown open and so didn't suck down the Sorensen and destroy everything.

Plugged everything back in switched the WFS servos back on. Looks good. Took before and after spectra.

In the plot:

GREEN: Open loop dark noise before changes
RED: Open loop bright (MC locked but MCWFS off)
BLUE: Closed loop, MC locked

BLACK: Dark noise after whitening
ORANGE:Closed loop after whitening

The cursor is at 16.25 Hz, the SOS bounce mode.

The I ran the new setMCWFSgains script which uses pzgain to set the UGFs of the 4 loops to 4.01 Hz.
We have in the past had problems with high WFS gains causing instabilities with the CM servo around 10-30 Hz. If this happens we should
just lower the gain by a factor of ~5.
Attachment 1: mcnoise.png
mcnoise.png
ELOG V3.1.3-